OpenStudy (anonymous):

Identify the period,x intercept and asymptones... F(x)=1/2tan(4x)

OpenStudy (anonymous):

The period would be the same as tan(4x). The asymptotes occur when you divide by a 0 (i.e. tan(4x)=0).

OpenStudy (anonymous):

im still confused... I really dont understand this

OpenStudy (anonymous):

Anything with just a trig function has an element of repetition in the function (i.e. the function looks the same if you move the function left or right by a certain amount. That amount is called the period.)

OpenStudy (anonymous):

Plot your function using google (it can graph things how) or fooplot. Asymptotes are basically when the function approaches some value, but never quite reaches that value. Here, you can never have a value x=pi/8 (ie tan(4x)=0, so you are dividing by 0).

OpenStudy (anonymous):

However, as you approach x=pi/8 from the negative side, the line curves up and almost touches the x=pi/2 line, and same from the positive side. But the positive side curves upwards instead of downwards

OpenStudy (anonymous):